OpenStudy (anonymous):

Kristen wants to hang three wedding pictures in the hallway of her home. She is choosing 3 pictures from a total of 10. If the order in which she hangs the pictures is not important, how many different ways can Kristen choose 3 pictures and hang them on the wall?

OpenStudy (anonymous):

\[\dbinom{10}{3}=\frac{10\times 9\times 8}{3\times 2}=10\times 3\times 4=120\]

OpenStudy (anonymous):

10 the order is NOT IMPORTANT so there are only 10

OpenStudy (anonymous):

123 456 789 012 345 678 901 234 567 890